6.27 yuan per S 62.37 rupee per S China 27 yuan India 150 rupee Compared to the U.S., prices in Switzerland are102.1__percent higher (circle one). If you took $50,000 to Switzerland it would feel like you had $ 24,742.27after purchasing power parity adjustment. Compared to the U.S., prices in India are you took S50,000 to India it would feel like you had S parity adjustment. percent higher/lower (circle one). If after purchasing power

Need help, please show work. Some correct answers are shown, all I need to know is how to work through the problem. Please show work.
Answer #1

Implied exchange rate = Price of Latte/Price of Latte in United States.

So, Implied exchange rate in Switzerland = 6.90/3.55 = 1.94 francs per $
Implied exchange rate in Sweden = 37.46/3.55 = 10.55 kronas per $
Implied exchange rate in China = 27/3.55 = 7.61 yuans per $
Implied exchange rate in India = 150/3.55 = 45.25 rupees per $

Compared to the US, dollar prices in Switzerland are 6.90 francs = 6.90/0.96 dollars = $7.19
So, percentage change in price in Switzerland = [(price in Switzerland-price in US)/price in US]*100 = [(7.19-3.55)/3.55]*100 = (3.64/3.55)*100 = 102.54 percent. As it is positive, so compared to the US, prices in Switzerland are higher.
If you took $50,000 to Switzerland, after purchasing power adjustment, $50,000 will equal, 50,000*(official exchange rate/implied exchange rate) = 50000*(0.96/1.94) = 48000/1.94 = $24,742.27

Compared to the US, dollar prices in India are 150 rupee = 150/62.37 dollars = $2.41
So, percentage change in price in India = [(price in India-price in US)/price in US]*100 = [(2.41-3.55)/3.55]*100 = (-1.14/3.55)*100 = -32.11 percent. As it is negative, so compared to the US, prices in India are lower.
If you took $50,000 to India, after purchasing power adjustment, $50,000 will equal, 50,000*(official exchange rate/implied exchange rate) = 50000*(62.37/45.25) = 3,118,500/45.25 = $68,917.13